# The Nonlinear Filter Model of Stream Cipher Redivivus

### v2

@@ 4 Concrete Proposals
The basic correlation attack [41] is applicable to the combiner generator model. Applying this attack to the filter generator model results in going through all the possible $2^{L}$ states of the LFSR. Since by our choice $L\geq 2\kappa$, the basic correlation attack does not defeat the $\kappa$-bit security level.
@@
Given $\kappa$, we obtained representative values of $L$ and $m$. The procedure we followed is to choose the value of $L$ to be the first prime number greater than $2\kappa$, and then for the chosen value of $L$, choose $m$ to be the least integer such that $\beta,\gamma>2^{\kappa}$. This ensures security against algebraic attacks considered above. Next, using $\kappa$ and the corresponding values of $L$ and $m$, we computed the maximum value of $B$ such that $B\leq\kappa$, $\alpha>2^{B}$, and (9), (10) and (11) hold. In each of the cases that we considered, it turns out that this maximum value of $B$ is in fact $\kappa$. So for the chosen values of $L$ and $m$, the corresponding $\mathcal{S}(L,m)$ ensures $\kappa$-bit security against the above analysed correlation attacks even when the adversary has access to $2^{\kappa}$ keystream bits. We note, however, that generating $2^{\kappa}$ keystream bits from a single key and IV pair is meaningless from a practical point of view. Instead, we set $B=64$, i.e. from a single key and IV pair at most $2^{64}$ keystream bits are to be generated. Table 1 shows the values of $\kappa$, $L$, $m$ and other parameters of the filtering function. We note the following points regarding the entries in Table 1.
@@ Appendix A Improved Degree and Algebraic Resistance
Suppose we set $\pi$ in the definition of ${\sf MM}_{2m}$ to be $m\mbox{-}{\sf HWBP}$ (along with choosing $h$ to be ${\sf Maj}_{m}$), and define ${\sf MM}_{2m+1}$ from ${\sf MM}_{2m}$ as in (4). This leads to an improvement of algebraic resistance. We conducted experiments to compute the AI and FAI of ${\sf MM}_{n}$ for $n$ up to 20. Our experiments suggest that $\lfloor n/3\rfloor\leq{\sf AI}({\sf MM}_{n})\leq 1+\lfloor n/3\rfloor$, and $\lfloor n/2\rfloor\leq{\sf FAI}({\sf MM}_{n})\leq 1+\lfloor n/2\rfloor$. (This is only a conjecture, and we do not have a proof.)
@@
${\sf MM}_{2m+1}$ is a balanced function having nonlinearity $2^{2m}-2^{m}$ and degree at most $m$. A construction of balanced functions on $12m+$ 1 variables which has nonlinearity $2^{2m}-2^{m}$ and the maximum possible degree $2m$ is given by Theorem 10(a) of [38]. Dobbertin [21] gave a general construction of balanced Boolean functions on an even number of variables. The proposal was to modify a normal bent function on $2m$ variables by inserting a balanced function on $m$ variables on the flat where the bent function is constant. This is a recursive procedure, where the recursion terminates with a balanced function on an odd number of variables and possessing the maximum possible nonlinearity. A simple modification of this recursive procedure is to require the recursion to terminate with a balanced function on an odd number of variables possessing the maximum possible degree and as high nonlinearity as possible. Then the resulting function on $2m$ variables is balanced and has maximum degree. The terminating function can be chosen to be the maximum degree balanced function on odd number of variables obtained from the construction given by Theorem 10(a) of [38]. The net result is that for any $n$ (either odd or even), it is possible to obtain balanced functions with maximum degree and very high nonlinearity. The procedure requires bent functions at several stages. One may instantiate these bent functions using ${\sf MM}_{2m}$ with $\pi$ chosen to be $m\mbox{-}{\sf HWBP}$ and $h$ chosen to be ${\sf Maj}_{m}$. We have implemented this construction and computed the algebraic immunities of the resulting functions. From our experiments for $n$ up to 20, we observed that (as above) the algebraic immunity is either $\lfloor n/3\rfloor$, or $1+\lfloor n/3\rfloor$, and the fast algebraic immunity is either $\lfloor n/2\rfloor$ or $1+\lfloor n/2\rfloor$.
@@
Even though the above modifications result in function with higher degrees and higher algebraic immunities, we did not opt for these functions. The reason is the implementation efficiency. The modifications make the functions more complex, which results in more gate requirements. In particular, the implementation of $m\mbox{-}{\sf HWBP}$ requires $O(m^{2})$ gates, whereas the implementation of ${\sf Maj}_{m}$ requires $O(m)$ gates. Since implementation efficiency is a primary concern, we choose the simpler option and compensate by increasing the number of variables.

### v3

@@ 4 Concrete Proposals
The basic correlation attack [53] is applicable to the combiner generator model. Applying this attack to the filter generator model results in going through all the possible $2^{L}$ states of the LFSR. Since by our choice $L\geq 2\kappa$, the basic correlation attack does not defeat the $\kappa$-bit security level. (An early correlation attack [52] on the nonlinear filter generator finds an equivalent representation of the stream cipher when the filtering function is not known; since we assume that the filtering function is known, this attack is not relevant to our context.)
@@
Concrete choices of $L$ and $m$.
Given $\kappa$, we obtained representative values of $L$ and $m$. The procedure we followed is to choose the value of $L$ to be the first prime number greater than $2\kappa$, and then for the chosen value of $L$, choose $m$ to be the least integer such that $\beta,\gamma>2^{\kappa}$. This ensures security against (fast) algebraic attacks considered above. The number of variables of the filtering function is $n=2m+1$. For each of the obtained values of $L$ and $n$, we checked whether (9) holds. In each case we found that (9) indeed holds. So the obtained values of $L$ and $m$ ensure security against state guessing attacks. Next, using $\kappa$ and the corresponding values of $L$ and $m$, we computed the maximum value of $B$ such that $B\leq\kappa$, $\alpha>2^{B}$, and (10), (11) and (12) hold. In each of the cases that we considered, it turns out that this maximum value of $B$ is in fact $\kappa$. So for the chosen values of $L$ and $m$, the corresponding $\mathcal{S}(L,m)$ ensures $\kappa$-bit security against the above analysed correlation attacks even when the adversary has access to $2^{\kappa}$ keystream bits. We note, however, that generating $2^{\kappa}$ keystream bits from a single key and IV pair is meaningless from a practical point of view. Instead, we set $B=64$, i.e. from a single key and IV pair at most $2^{64}$ keystream bits are to be generated. Table 1 shows the values of $\kappa$, $L$, $m$ and other parameters of the filtering function. We note the following points regarding the entries in Table 1.
@@ Appendix A Improved Degree and Algebraic Resistance
Dobbertin [25] had proposed a method for modifying normal bent functions to construct highly nonlinear balanced Boolean functions on an even number of variables. The method is recursive and requires balanced Boolean functions on an odd number of variables. Suppose that for odd $n$, ${\sf Bal}_{n}$ is a balanced function on $n$ variables. Dobbertin’s recursive method applied to the Maiorana-McFarland bent functions is the following.
Suppose $\pi$ is the permutation used to define ${\sf MM}_{2m}$. Let $\mathbf{a}=(a_{1},\ldots,a_{m})=\pi^{-1}(\mathbf{0}_{m})$. We define $\mathbf{1}_{\mathbf{a}}(\mathbf{X})=(1\oplus a_{1}\oplus X_{1})\cdots(1\oplus a_{1}\oplus X_{m})$. If $\pi$ is a bit permutation, then $\mathbf{a}=\mathbf{0}_{m}$ and $\mathbf{1}_{\mathbf{a}}(\mathbf{X})$ is simply $(1\oplus X_{1})\cdots(1\oplus X_{m})$. For $m\geq 1$, ${\sf Bal}_{2m}$ is defined as follows.
${\sf Bal}_{2m}(\mathbf{X},\mathbf{Y})$ $=$ ${\sf MM}_{2m}(\mathbf{X},\mathbf{Y})\oplus\mathbf{1}_{\mathbf{a}}(\mathbf{X}){\sf Bal}_{m}(\mathbf{Y}).$ (15)
The recursive definition in (15) terminates when $m$ is odd. It is easy to see that for $m\geq 1$, ${\sf deg}({\sf Bal}_{2m})=m+{\sf deg}({\sf Bal}_{m})$, and ${\sf nl}({\sf Bal}_{2m})=2^{2m-1}-2^{m}+{\sf nl}({\sf Bal}_{m})$. The following result provides a lower bound on the algebraic immunity of ${\sf Bal}_{2m}$.
Theorem 2.
${\sf AI}({\sf Bal}_{2m})$$\geq$ ${\sf wt}(\bm{\omega}^{\star})+{\sf AI}\big{(}\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\oplus{\sf Bal}_{m}(\bm{\omega}^{\star})\mathbf{1}_{\mathbf{a}}(\mathbf{X})\big{)}.$
Proof.
Suppose $g(\mathbf{X},\mathbf{Y})$ is an annihilator for ${\sf Bal}_{2m}(\mathbf{X},\mathbf{Y})$, where $g(\mathbf{X},\mathbf{Y})\allowbreak=\allowbreak\bigoplus_{\bm{\omega}\in\mathbb{F}_{2}^{m}}\allowbreak\mathbf{Y}^{\bm{\omega}}g_{\bm{\omega}}(\mathbf{X})$ for some functions $g_{\bm{\omega}}(\mathbf{X})$. Also, we write ${\sf Bal}_{m}(\mathbf{Y})=\bigoplus_{\bm{\omega}\in\mathbb{F}_{2}^{m}}b_{\bm{\omega}}\mathbf{Y}^{\bm{\omega}}$, where $b_{\bm{\omega}}\in\mathbb{F}_{2}$. We have
00 0 $=$ $g(\mathbf{X},\mathbf{Y}){\sf Bal}_{2m}(\mathbf{X},\mathbf{Y})$ (16)
$=$ $\left(\bigoplus_{\bm{\omega}\in\mathbb{F}_{2}^{m}}\mathbf{Y}^{\bm{\omega}}g_{\bm{\omega}}(\mathbf{X})\right)$
$\left(\pi_{1}(\mathbf{X})Y_{1}\oplus\cdots\oplus\pi_{m}(\mathbf{X})Y_{m}\oplus h(\mathbf{X})\oplus\mathbf{1}_{\mathbf{a}}(\mathbf{X})\bigoplus_{\bm{\omega}\in\mathbb{F}_{2}^{m}}b_{\bm{\omega}}\mathbf{Y}^{\bm{\omega}}\right).$
Since the right hand side of (16) is equal to 0, all coefficients of $\mathbf{Y}^{\bm{\omega}}$ in the expansion on the right hand side of (16) must be equal to 0. Since $g(\mathbf{X},\mathbf{Y})\neq 0$, let $w\geq 0$ be the minimum integer such that there is an $\bm{\omega}^{\star}$ with ${\sf wt}(\bm{\omega}^{\star})=w$ and $g_{\bm{\omega}^{\star}}(\mathbf{X})\neq 0$. In (16), equating the coefficient of $\mathbf{Y}^{\bm{\omega}^{\star}}$ to 00, we have
00 0 $=$ $g_{\bm{\omega}^{\star}}(\mathbf{X})\left(h(\mathbf{X})\oplus\left(\bigoplus_{i\in{\sf supp}(\bm{\omega}^{\star}))}\pi_{i}(\mathbf{X})\right)\oplus\left(\mathbf{1}_{\mathbf{a}}(\mathbf{X})\bigoplus_{\bm{\omega}\leq\bm{\omega}^{\star}}b_{\bm{\omega}}\right)\right)$
$=$ $g_{\bm{\omega}^{\star}}(\mathbf{X})\big{(}\langle\bm{\omega},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\oplus{\sf Bal}_{m}(\bm{\omega}^{\star})\mathbf{1}_{\mathbf{a}}(\mathbf{X})\big{)}.$
Here we have used ${\sf Bal}_{m}(\bm{\omega}^{\star})=\bigoplus_{\bm{\omega}\leq\bm{\omega}^{\star}}b_{\bm{\omega}}$ (see (1)). So $g_{\bm{\omega}^{\star}}(\mathbf{X})$ is an annihilator for $\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\oplus{\sf Bal}_{m}(\bm{\omega}^{\star})\mathbf{1}_{\mathbf{a}}(\mathbf{X})$. Consequently, ${\sf deg}(g)\geq{\sf wt}(\bm{\omega}^{\star})+{\sf deg}(g_{\bm{\omega}^{\star}})\geq{\sf wt}(\bm{\omega}^{\star})+{\sf AI}(\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\oplus{\sf Bal}_{m}(\bm{\omega}^{\star})\mathbf{1}_{\mathbf{a}}(\mathbf{X}))$.
If, on the other hand, $g(\mathbf{X},\mathbf{Y})$ is an annihilator for $1\oplus{\sf MM}_{2m}(\mathbf{X},\mathbf{Y})$, then a similar argument shows that $g_{\bm{\omega}^{\star}}(\mathbf{X})$ is an annihilator for $\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus 1\oplus h(\mathbf{X})\oplus{\sf Bal}_{m}(\bm{\omega}^{\star})\mathbf{1}_{\mathbf{a}}(\mathbf{X})$, and again we have ${\sf deg}(g)\geq{\sf wt}(\bm{\omega}^{\star})+{\sf AI}(\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\oplus{\sf Bal}_{m}(\bm{\omega}^{\star})\mathbf{1}_{\mathbf{a}}(\mathbf{X}))$.
The lower bound on ${\sf AI}({\sf MM}_{2m})$ is in terms of the AI of the function $f=\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})$, while the lower bound on ${\sf AI}({\sf Bal}_{2m})$ is in terms of the AI of the function $g=\langle\bm{\omega}^{\star},\pi(\mathbf{X})\rangle\oplus h(\mathbf{X})\oplus{\sf Bal}_{m}(\bm{\omega}^{\star})\mathbf{1}_{\mathbf{a}}(\mathbf{X})$. Note that $f$ and $g$ differ in at most one bit, which shows that the AI of ${\sf AI}({\sf MM}_{2m})$ and ${\sf AI}({\sf Bal}_{2m})$ are quite close.
The computation of ${\sf MM}_{2m}$ requires the computation of several Maiorana-McFarland functions on a smaller number of variables, terminating with the computation of a balanced function on an odd number of variables. If we instantiate the function $h$ with the majority function, then the computation of ${\sf MM}_{2m}$ requires several inner product and majority function computations on smaller number of variables. These push up the circuit size of ${\sf Bal}_{2m}$ in comparison to that of ${\sf MM}_{2m}$, but do not lead to any substantial gain in the AI.
@@
Suppose we set $\pi$ in the definition of ${\sf MM}_{2m}$ to be $m\mbox{-}{\sf HWBP}$ (along with choosing $h$ to be ${\sf Maj}_{m}$), and define ${\sf MM}_{2m+1}$ from ${\sf MM}_{2m}$ as in (4). This leads to an improvement of algebraic resistance. We conducted experiments to compute the AI and FAI of ${\sf MM}_{n}$ for $n$ up to 20. The results are shown in Tables 5 and 5, and suggest that $\lfloor n/3\rfloor\leq{\sf AI}({\sf MM}_{n})\leq 1+\lfloor n/3\rfloor$, and $\lfloor n/2\rfloor\leq{\sf FAI}({\sf MM}_{n})\leq 1+\lfloor n/2\rfloor$.
@@
A construction of balanced functions on $12m+$ 1 variables which has nonlinearity $2^{2m}-2^{m}$ and the maximum possible degree $2m$ is given by Theorem 10(a) of [49]. This construction requires bent functions on smaller number of variables. Suppose we choose these bent functions to ${\sf MM}_{2m}$ with $\pi$ built from ${\sf HWBP}$ and $h$ built from the majority function on appropriate number of variables. For odd $n$, let us define ${\sf Bal}_{n}$ to be such a function. Suppose that the terminating function in (15) is chosen to be ${\sf Bal}_{n}$. The net result is that for any $n$ (either odd or even), it is possible to obtain balanced functions with maximum degree and very high nonlinearity. We have implemented this construction and computed the algebraic immunities and fast algebraic immunities of the resulting functions. The results of our experiments for $n$ up to 20 are shown in Tables 5 and 5, and again we observe that the algebraic immunity is either $\lfloor n/3\rfloor$, or $1+\lfloor n/3\rfloor$, and the fast algebraic immunity is either $\lfloor n/2\rfloor$ or $1+\lfloor n/2\rfloor$.
$n$ ${\sf AI}({\sf Maj}_{\lceil n/2\rceil})$ ${\sf AI}({\sf HWB}_{\lceil n/2\rceil})$ ${\sf AI}({\sf MM}_{n})$ ${\sf AI}({\sf Bal}_{n})$
4 1 1 2 2
5 1 1 2 2
6 2 2 3 3
7 2 2 3 3
8 2 2 3 3
9 2 2 3 4
10 3 2 4 4
11 3 2 4 4
12 3 3 5 5
13 3 3 5 5
14 4 3 5 5
15 4 3 6 6
16 4 4 5 5
17 4 4 6 6
18 5 4 7 7
19 5 4 7 7
20 5 4 7 7
Table 4: Algebraic immunities of ${\sf MM}$ and ${\sf Bal}$, where $\pi$ is chosen to be $m\mbox{-}{\sf HWBP}$ and $h$ is chosen to be ${\sf Maj}_{m}$.
$n$ FAA-profile of ${\sf MM}_{n}$ ${\sf FAI}({\sf MM}_{n})$ FAA-profile of ${\sf Bal}_{n}$ ${\sf FAI}({\sf Bal}_{n})$
4 (1,1) 3 (1,1) 3
5 (1,1) 3 (1,2) 4
6 (1,2), (2,2) 4 (1,2), (2,2) 4
7 (1,2), (2,2) 4 (1,2), (2,2) 4
8 (1,3), (2,2) 5 (1,3), (2,2) 5
9 (1,3), (2,2) 5 (1,3), (2,3), (3,3) 5
10 (1,4), (2,3), (3,3) 6 (1,4), (2,3), (3,3) 6
11 (1,4), (2,3), (3,3) 6 (1,4), (2,4), (3,3) 6
12 (1,4), (2,4), (3,4), (4,4) 6 (1,5), (2,4), (3,4), (4,4) 7
13 (1,4), (2,4), (3,4), (4,4) 6 (1,5), (2,4), (3,4), (4,4) 7
14 (1,6), (2,5), (3,5), (4,5) 8 (1,6), (2,6), (3,5), (4,5) 8
15 (1,6), (2,5), (3,5), (4,5), (5,5) 8 (1,6), (2,5), (3,5), (4,5), (5,5) 8
16 (1,7), (2,6), (3,5), (4,5) 9 (1,7), (2,6), (3,5), (4,5) 9
17 (1,7), (2,6), (3,5), (4,5), (5,5) 9 (1,7), (2,6), (3,5), (4,5), (5,5) 9
18 (1,8), (2,7), (3,7), (4,6) (1,8), (2,7), (3,7), (4,6)
19 (1,8), (2,7), (3,7) (1,8), (2,7), (3,7)
20 (1,8), (2,8), (3,8) (1,9), (2,8), (3,8)
Table 5: (Partial) FAA-profiles and FAI of ${\sf MM}$ and ${\sf Bal}$, where $\pi$ is chosen to be $m\mbox{-}{\sf HWBP}$ and $h$ is chosen to be ${\sf Maj}_{m}$.
@@
Even though ${\sf Bal}_{n}$ has higher degree and higher algebraic immunities, we did not opt for ${\sf Bal}_{n}$. The reason is the implementation efficiency. The modifications make the functions more complex, which results in more gate requirements. In particular, the implementation of $m\mbox{-}{\sf HWBP}$ requires $O(m^{2})$ gates, whereas the implementation of ${\sf Maj}_{m}$ requires $O(m)$ gates. Since implementation efficiency is a primary concern, we choose the simpler option and compensate by increasing the number of variables.